# Isometries of spacetimes without observer horizons

### v2

@@ 3.2. Examples
In the above proof, compactness of $\Sigma$ is essential. Indeed, Minkowski spacetime $\mathbb{R}^{1,n}$ is also a product spacetime, but its isometry group, the Poincaré group, includes boosts (which mix space and time), and hence is larger than $\mathbb{R}\times\mathsf{O}(n)$. Note also that product spacetimes with compact slices always satisfy the NOH. An example where the NOH is violated and the conclusion of Corollary 1.3 fails, is given by de Sitter spacetime $\mathsf{dS}^{4}$. On the one hand, $\mathsf{dS}^{4}$ can be represented as the warped product $\mathsf{dS}^{4}=\mathbb{R}^{1,0}\times_{\cosh^{2}}S^{3}$, implying that is is globally hyperbolic with compact Cauchy surfaces [4, pp. 183–184]. On the other hand, $\mathsf{dS}^{4}$ can also be represented as the pseudosphere in Minkowski spacetime $\mathbb{R}^{1,4}$. It follows that its isometry group is the Lorentz group $\mathsf{O}^{\uparrow}(1,4)$ [29, Prop. 9.8], which, again because it contains boosts, cannot be written in the form $\mathbb{R}\ltimes N$ with $N$ compact. While the obvious examples of spacetimes with isometry group $\mathbb{R}\ltimes N$ are the ones that admit a complete timelike Killing vector field (such as $\partial_{t}$ in the case of a product spacetime), the following example shows that they are not the only ones.
@@
Note that the level sets of the $t$-coordinate are compact Cauchy surfaces, and that the periodic dependence of $g$ in $t$ implies that $(M,g)$ satisfies the NOH [16, Thm. 1.6]. We are going to show that the subgroup of $\operatorname{Isom}^{\uparrow}(M,g)$ preserving an invariant time function is
@@
is a Riemannian isometry, which is impossible, because the circumference of the smaller circle in $(\Sigma_{t},h_{t})$ is an invariant, and in this case, they do not match.
@@
defines a $G$-invariant Lorentzian metric on $M$. As in the previous example, the level sets $\Sigma_{t}$ are compact Cauchy surfaces, and the periodic dependence of $g$ in $t$ (due to $G$-invariance) implies that $(M,g)$ satisfies the NOH [16, Thm. 1.6].
@@ 4.1. Uniform Lipschitzianity of isometries
On the one hand, because the image of each $\gamma_{i}$ lies in $U$, we have that the image of each $\sigma_{i}$ lies in $K^{\prime}$. On the other hand, if we reparametrize the $\sigma_{i}$ with respect to $h$-arclength, then by the Lorentzian limit curve theorem [4, Prop. 3.31], $\sigma_{i}$ converges (up to a subsequence) to a limit causal geodesic $\sigma$ with image contained in $K^{\prime}$. But up to reparametrization, $\sigma$ can be nothing else than the null geodesic
@@ 4.2. Spacetimes with a regular cosmological time function
It is a consequence of the above proof (combined with [12, Cor. 1]) that the level sets of the cosmological time function for small values (smaller than $\varepsilon$) are Cauchy surfaces. This had previously been proven in [14] for all level sets, under the additional assumption that the spacetime is future timelike geodesically complete. Theorem 4.2 also applies to the regular cosmological volume functions introduced in [15, Sec. 6], as these are also, by construction, invariant under isometries.

### v3

@@ 3.2. Examples
In the above proof, compactness of $\Sigma$ is essential. Indeed, Minkowski spacetime $\mathbb{R}^{1,n}$ is also a product spacetime, but its isometry group, the Poincaré group, includes boosts (which mix space and time), and hence is larger than $\mathbb{R}\times\mathsf{O}(n)$. Note also that product spacetimes with compact slices always satisfy the NOH. An example where the NOH is violated and the conclusion of Corollary 1.3 fails, is given by de Sitter spacetime $\mathsf{dS}^{4}$. On the one hand, $\mathsf{dS}^{4}$ can be represented as the warped product $\mathsf{dS}^{4}=\mathbb{R}^{1,0}\times_{\cosh^{2}}S^{3}$, implying that it is globally hyperbolic with compact Cauchy surfaces [5, pp. 183–184]. On the other hand, $\mathsf{dS}^{4}$ can also be represented as the pseudosphere in Minkowski spacetime $\mathbb{R}^{1,4}$. It follows that its isometry group is the Lorentz group $\mathsf{O}^{\uparrow}(1,4)$ [32, Prop. 9.8], which, again because it contains boosts, cannot be written in the form $\mathbb{R}\ltimes N$ with $N$ compact. While the obvious examples of spacetimes with isometry group $\mathbb{R}\ltimes N$ are the ones that admit a complete timelike Killing vector field (such as $\partial_{t}$ in the case of a product spacetime), the following example shows that they are not the only ones.
@@
Note that the level sets of the $t$-coordinate are compact Cauchy surfaces, and that the periodic dependence of $g$ in $t$ implies that $(M,g)$ satisfies the NOH [19, Thm. 1.6]. We are going to show that the subgroup of $\operatorname{Isom}^{\uparrow}(M,g)$ preserving an invariant time function is
@@
is a Riemannian isometry, which is impossible, because the smaller radius of each torus $(\Sigma_{t},h_{t})$ is an invariant, and in this case, they do not match.
@@
defines a $G$-invariant Lorentzian metric on $M$. As in the previous example, the level sets $\Sigma_{t}$ are compact Cauchy surfaces, and the periodic dependence of $g$ in $t$ (due to $G$-invariance) implies that $(M,g)$ satisfies the NOH [19, Thm. 1.6].
@@ 4.1. Uniform Lipschitzianity of isometries
On the one hand, because the image of each $\gamma_{i}$ lies in $U$, we have that the image of each $\sigma_{i}$ lies in $K^{\prime}$. On the other hand, if we reparametrize the $\sigma_{i}$ with respect to $h$-arclength, then by the Lorentzian limit curve theorem [5, Prop. 3.31], $\sigma_{i}$ converges (up to a subsequence) to a limit causal geodesic $\sigma$ with image contained in $K^{\prime}$. But up to reparametrization, $\sigma$ can be nothing else than the null geodesic
@@ 4.2. Spacetimes with a regular cosmological time function
It is a consequence of the above proof (combined with [16, Cor. 1]) that the level sets of the cosmological time function for small values (smaller than $\varepsilon$) are Cauchy surfaces. This had previously been proven in [14] for all level sets, under the additional assumption that the spacetime is future timelike geodesically complete. Theorem 4.2 also applies to the regular cosmological volume functions introduced in [18, Sec. 6], as these are also, by construction, invariant under isometries.